OpenStudy (boxman61):

An iron ball with a mass of 128.45 g and a diameter of 3.92 cm is placed in a graduated cylinder containing 40 mL of water. The water level rises to 72 mL. Find the density of the ball and compare this to the density of iron. Explain what is going on.

OpenStudy (boxman61):

I'm not to sure about this one... D=m/v M=128.45 V= 72mL-40mL= 32mL D= 128.45/32 = 4.02g/mL

OpenStudy (boxman61):

Density of Iron (Fe) is 7.874g/cm^3

OpenStudy (boxman61):

as 1mL is equal to 1cm^3 that gives me a density difference of 3.864g/cm^3 I'm not sure why other than the diameter of the ball but that shouldn't effect irons density unless surface area has something to do with it...

OpenStudy (boxman61):

As I thought, surface area does not effect density

OpenStudy (agent0smith):

It's not iron.

TheSmartOne (thesmartone):

Your work looks good. The diameter is useless information. Also, if the density is not the same density as iron that means they misidentified the ball as iron.

OpenStudy (boxman61):

I'll have to email my instructor about this one and find out whats going on.. thank you both

OpenStudy (agent0smith):

"Explain what is going on." can't you just say it's not iron? It's an alloy probably.

OpenStudy (boxman61):

I will present that as my conclusion and see if its correct... I don't know how it could be iron